how do you graph 3x+4y=-1, and how do you get x on one side to be an equation?
Answer by waynest(281)

3X + 4Y = -1
3X - 3X + 4Y = -3X -1
4Y = -3X -1
4Y/4 = -3X/4 -1/4
Y = -3/4X - 1/4